Case studies Availability and Consequences of Credit Purchases Mr.Krishna is willing to buy a Motor Lorry of R 1000000 from the Mayur Motor Company by paying cash of Rs 2,00,000. He want to pay the balance through instalments of Rs 1.60.000 rachat an interest of 5% on 31 December every year for 5 years Krishna wants to know about the different financial services provided by the Mayur Motor Company. As u Finance Manager of Mayur Motor Company, you are required to suggest Mr.

Krishna Regarding the types of Credit Purchase be can uvail from your concern

What will be the total interest borne by Mr. Krishna for the credit purchase made by him?

What consequences will the purchaser face

li the non-payment of installment Advise
1
Expert's answer
2020-03-03T09:34:57-0500

We calculate the balance of payment

1 000 000-2 00 000=800 000

Installment plan for 5 years:

800 000/5=160 000

Сalculate the total loan debt:


Date Payment Interest Loan body Balance


01.03.2020 0 0 0 800 000.00

01.03.2021 184 779.84 40 000.00 144 779.84 655 220.16

01.03.2022 184 779.84 32 761.01 152 018.83 503 201.33

01.03.2023 184 779.84 25 160.07 159 619.77 343 581.56

01.03.2024 184 779.84 17 179.08 167 600.76 175 980.80

01.03.2025 184 779.84 8 799.04 175 980.80 0

Total 923 899.20 123 899.20 800 000.00 0


The financial manager must offer a loan

Total percentage: 923 899.20/800 000=1.1549 or 15.49%

In case of non-payment by installments there will be penalties